OpenStudy (anonymous):

You need 1 1/4 cups of sugar to make 20 cookies. How many cups of sugar will you need to make 6 cookies?

OpenStudy (anonymous):

create a proportion sugar/cookies 1.25/20=x/14 cross multiply and solve for x 14(1.25)=20x divide by 20 .875 or 7/8 = x

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }\] of a cup.